OpenStudy (jozelynw):

How do you make a exponential graph that represents f(x)=100(1+0.7)^x ?

OpenStudy (jozelynw):

@mathmate

OpenStudy (jozelynw):

@peachpi

OpenStudy (jozelynw):

@mathmate @Thesmarterone

OpenStudy (mathmate):

Hint:  Here you have a=100, b=1.07 just need a few more points to give the shape of the graph.

OpenStudy (jozelynw):

you don't need the points of x and y> @mathmate

OpenStudy (mathmate):

The graph that I drew free-hand is sufficient for most purposes. Some teachers require at least one more point so that the graph defines completely the function.

OpenStudy (mathmate):

The graph that I drew free-hand is sufficient for most purposes. Some teachers require at least one more point so that the graph defines completely the function.

OpenStudy (jozelynw):

@mathmate what point? like f(1)=170 ,a point like this.

OpenStudy (mathmate):

Looked at it again, the graph is actually \(y=100(1.7^x)\), my bad! Except for b=1.7 (and not 1.07), the graph holds. Typically in a graph of an exponential function \(f(x)=a(b^x)\) f(0)=a, (because x=0, f(0)=\(ab^0=a, because b^0=1\) This point was indicated on the graph. f(1)=ab (because x=1, f(1)=\(ab^1=ab\)

OpenStudy (mathmate):

Yes, one more point like f(1)=170 defines the function completely, since a and b will then be defined.